OpenStudy (anonymous):

A culture started with 5,000 bacteria. After 3 hours, it grew to 6,500 bacteria. How many bacteria will be present after 13 hours? Use the formula P=Ae^kt

OpenStudy (anonymous):

5000^k*3=6500-5000=1500 5000^k=500 5000^k*(3+13)=500*16=8000 totally 5000+8000=13000 bacteria

OpenStudy (anonymous):

er... @marsss \[P = Ae^{kt}\]

OpenStudy (anonymous):

Have you learnt how to find derivatives yet?

OpenStudy (anonymous):

@completeidiot i tried what you say but it couldn't be solved.

OpenStudy (anonymous):

I HAVEN'T SLEPT IN TWO AND AHALF DAYS I CAN'T DO MATH, DON'T JUDGE ME.

OpenStudy (anonymous):

nonsense starting with the initial state at time t=0, there is a total of 5000 cells

OpenStudy (anonymous):

@AnImEfReaK are you a judge?

OpenStudy (anonymous):

No just helping.

OpenStudy (anonymous):

That's what I've been taught to do here...

OpenStudy (anonymous):

using the equation \[P=Ae^{kt}\] P = the total amount of bacteria at time t t= time lets say in hours A and k are constants which you will need to solve for and are relatively easy to solve for given the information now given the initial information, we can substitute the values in to get \[5000 = Ae^{k * 0}\] now solve for A

OpenStudy (anonymous):

now, note that anything to the zeroth power is equal to 1, and thus the e^kt part "cancels out" or becomes one, and you will be left with A = 5000

OpenStudy (anonymous):

@completeidiot you're really an idiot

OpenStudy (anonymous):

now, given the 2nd set of data after 3 hours, the total number of bacteria within the container is 6500 again using the equation \[P=Ae^{kt}\] since we already solve for A and we know now that P= 6500 and t =3 we substitute those in to get \[6500 = 5000 e^{k*3}\] now at this point, you want to solve for k

OpenStudy (anonymous):

@completeidiot you're wrong, can't you see?

OpenStudy (anonymous):

THIS IS A WASTE OF TIME.

OpenStudy (anonymous):

@marsss please kindly point out my mistake instead of concluding that i am wrong, you would be much more productive that way

OpenStudy (anonymous):

@ughwhy_ dude, is my answer right?

OpenStudy (anonymous):

anyways continuing on \[6500= 5000 e^{k*3}\] dividing both sides by 5000 \[\frac{6500}{5000} = e ^{k *3}\] taking the natural log of both sides \[\ln{ \frac{6500}{5000}}= 3*k\] at this point, divide both sides by 3 to isolate k

OpenStudy (anonymous):

since you now know k and you also know A you can substitute these values back into the given equation \[P=Ae^{kt}\] and then to determine the number of bacteria after 13 hours, make the substitution t=13 and solve for P

OpenStudy (anonymous):

@completeidiot congratulations. you solved that like a genius

OpenStudy (anonymous):

@UnkleRhaukus @dumbcow @.Sam. please check my explanation/work and marsss' work

OpenStudy (unklerhaukus):

@completeidiot +1